Consider the following functions.

f={(2,5),(3,−2),(0,0)}
and

g={(−2,−5),(3,1)}

Find (f+g)(3) =
Answer #1

Answer: -1

Proof:

(f+g)(x) = f(x) + g(x)

f(3) + g(3) = (-2) + (1) = -1

answered by: Guest
Answer #2

the answer is really simple it is the cirucumfrence of my foreskin divided by the width of my 8===D - - 

   

source: wikepedia
answered by: g_floyd6969